# A contramodule generalization of Neeman’s flat and projective module theorem

### v1

@@ 6. Contraacyclic Complexes of Flat Contramodules are Acyclic
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ of complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. The $\aleph_{1}$-presentable objects of this category are precisely all the complexes of countably presentable flat left $\mathfrak{R}$-contramodules.
@@
This is [26, Proposition 10.2]. The assertion follows from Propositions 4.6(a) and 1.2(b). ∎
@@
Since the class of all acyclic complexes in an exact category is closed under kernels of termwise admissible epimorphisms and cokernels of termwise admissible monomorphisms, the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic. On the other hand, in view of Lemma 2.4, the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. So we can consider $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ instead of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and assume that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. Our aim is to prove that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is absolutely acyclic. What we have achieved here is that we know additionally that all terms of the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are quotient contramodules of cotorsion contramodules.
@@
It is easy to represent $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ as a quotient complex of a complex of (countably presentable) projective $\mathfrak{R}$-contramodules $\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. So we have a short exact sequence of complexes of $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$. By Proposition 4.7,  $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is also a complex of projective $\mathfrak{R}$-contramodules (in fact, of countably presentable projective $\mathfrak{R}$-contramodules, by [26, Lemma 11.1]). Put $\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}}=\operatorname{Tot}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$. Then the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic by Corollary 6.5. The complexes of abelian groups $\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ and $\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic since the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ by assumption. Thus the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic, too.
@@
7. Projective and Flat Constructions of Contraderived Category
Now we can formulate the theorem promised in the introduction, summarizing our results generalizing Neeman’s [18, Theorem 8.6].

### v3

@@ 6. Contraacyclic Complexes of Flat Contramodules are Acyclic
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ of complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. A complex of flat $\mathfrak{R}$-contramodules $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is $\aleph_{1}$-presentable as an object of $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ if and only if all the $\mathfrak{R}$-contramodules $\mathfrak{H}^{n}$,  $n\in\mathbb{Z}$, are countably presentable.
@@
This is [34, Proposition 10.2]. The assertion follows from Propositions 4.6(a) and 1.2(b) (with $\kappa=\aleph_{1}$ and $\lambda=\aleph_{0}$). ∎
@@
Since the class of all acyclic complexes in an exact category is closed under kernels of termwise admissible epimorphisms and cokernels of termwise admissible monomorphisms, the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic. On the other hand, in view of Lemma 2.5(a), the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic if and only if the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. So we can consider $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ instead of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and assume that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic. Our aim is to prove that $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic. What we have achieved here is that we know additionally that all the terms of the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are quotient contramodules of cotorsion contramodules.
@@
It is easy to represent $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ as a quotient complex of a complex of (countably generated) projective $\mathfrak{R}$-contramodules $\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$. So we have a short exact sequence of complexes of $\mathfrak{R}$-contramodules $0\longrightarrow\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$. By Proposition 4.7,  $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is also a complex of projective $\mathfrak{R}$-contramodules (in fact, of countably generated projective $\mathfrak{R}$-contramodules, by [34, Lemma 11.1]). Put $\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}}=\operatorname{Tot}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\to\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$. Then the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic by Corollary 6.5. The complexes of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ and $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic, since the complex $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is contraacyclic in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ by assumption. Thus the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic, too.
@@
7. Cotorsion Pairs
This section continues the discussion of category-theoretic background that was started in Section 2. Given an exact category $\mathsf{K}$, a pair of classes of objects $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ is said to be a cotorsion pair if $\mathsf{C}=\mathsf{F}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$.
For any class of objects $\mathsf{S}\subset\mathsf{K}$, the pair of classes $\mathsf{C}=\mathsf{S}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$ is a cotorsion pair in $\mathsf{K}$. The cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be generated by the class of objects $\mathsf{S}$.
Dually to the definition in Section 2, a class of objects $\mathsf{C}\subset\mathsf{K}$ is said to be cogenerating if for every object $K\in\mathsf{K}$ there exists an object $C\in\mathsf{C}$ together with an admissible monomorphism $K\longrightarrow C$ in $\mathsf{K}$.
Let $(\mathsf{F},\mathsf{C})$ be a cotorsion pair in $\mathsf{K}$ such that the class $\mathsf{F}$ is generating and the class $\mathsf{C}$ is cogenerating in $\mathsf{K}$. Then the cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be hereditary if any one of the following equivalent conditions holds:
  1. (1)
    the class $\mathsf{F}$ is closed under kernels of admissible epimorphisms in $\mathsf{K}$;
  2. (2)
    the class $\mathsf{C}$ is closed under cokernels of admissible monomorphisms in $\mathsf{K}$;
  3. (3)
    $\operatorname{Ext}^{2}_{\mathsf{K}}(F,C)=0$ for all $F\in\mathsf{F}$ and $C\in\mathsf{C}$;
  4. (4)
    $\operatorname{Ext}^{n}_{\mathsf{K}}(F,C)=0$ for all $F\in\mathsf{F}$,  $C\in\mathsf{C}$, and $n\geq 1$.
The nontrivial implications are (1) $\Longrightarrow$ (4) and (2) $\Longrightarrow$ (4); they hold by Lemma 2.1 and its dual version.
A cotorsion pair $(\mathsf{F},\mathsf{C})$ is $\mathsf{K}$ is said to be complete if for every object $K\in\mathsf{K}$ there exist (admissible) short exact sequences
[…]
in $\mathsf{K}$ with objects $F$, $F^{\prime}\in\mathsf{F}$ and $C$, $C^{\prime}\in\mathsf{C}$. The short exact sequence (1) is said to be a special precover sequence. The short exact sequence (2) is said to be a special preenvelope sequence.
The following result is known classically as the Eklof–Trlifaj theorem [12, Theorems 2 and 10]. Here, given a class of objects $\mathsf{F}\subset\mathsf{K}$, we denote by $\mathsf{F}^{\oplus}\subset\mathsf{K}$ the class of all direct summands of the objects from $\mathsf{F}$. The notation $\mathsf{Fil}(\mathsf{S})\subset\mathsf{K}$ was introduced in Section 2.
@@ 7. Cotorsion Pairs
Let $\mathsf{B}$ be a locally presentable abelian category and $\mathsf{S}\subset\mathsf{B}$ be a set of objects. Consider the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{B}$ generated by $\mathsf{S}$. In this context:
(a) if the class $\mathsf{F}$ is generating in $\mathsf{B}$ and the class $\mathsf{C}$ is cogenerating in $\mathsf{B}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ is complete;
(b) if the class $\mathsf{Fil}(\mathsf{S})$ is generating in $\mathsf{B}$, then $\mathsf{F}=\mathsf{Fil}(\mathsf{S})^{\oplus}$.
Proof.
Part (a) is [38, Corollary 3.6]. Part (b) is [38, Theorem 4.8(a,d)]. See also [40, Theorems 3.3 and 3.4]. ∎
The following result is closely related to Theorem 2.8.
Theorem 7.2.
Let $\mathsf{B}$ be a locally presentable abelian category with enough projective objects. Then the pair of classes of complexes of projective objects $\mathsf{F}=\mathsf{Com}(\mathsf{B}_{\mathsf{proj}})$ and contraacyclic complexes $\mathsf{C}=\mathsf{Ac}^{\mathsf{bctr}}(\mathsf{B})$ is a hereditary complete cotorsion pair $(\mathsf{F},\mathsf{C})$ in the abelian category of complexes $\mathsf{Com}(\mathsf{B})$.
Proof.
This is [40, Theorem 7.3]. For a generalization, see [41, Theorem 6.16]. ∎
Let $(\mathsf{F},\mathsf{C})$ be a complete cotorsion pair in $\mathsf{K}$ and $\mathsf{E}\subset\mathsf{K}$ be a full subcategory closed under extensions. We will say that the cotorsion pair $(\mathsf{F},\mathsf{C})$ restricts to a complete cotorsion pair in $\mathsf{E}$ if the pair of classes $\mathsf{E}\cap\mathsf{F}$ and $\mathsf{E}\cap\mathsf{C}$ is a complete cotorsion pair in (the inherited exact structure on) $\mathsf{E}$.
Lemma 7.3.
Let $\mathsf{K}$ be an exact category and $\mathsf{E}\subset\mathsf{K}$ be a full subcategory closed under extensions. Let $(\mathsf{F},\mathsf{C})$ be a complete cotorsion pair in $\mathsf{K}$. In this setting:
(a) if $\mathsf{E}$ is closed under kernels of admissible epimorphisms in $\mathsf{K}$ and $\mathsf{F}\subset\mathsf{E}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ restricts to a complete cotorsion pair $(\mathsf{F},\,\mathsf{E}\cap\mathsf{C})$ in $\mathsf{E}$;
(b) if $\mathsf{E}$ is closed under cokernels of admissible monomorphisms in $\mathsf{K}$ and $\mathsf{C}\subset\mathsf{E}$, then the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ restricts to a complete cotorsion pair $(\mathsf{E}\cap\mathsf{F},\,\mathsf{C})$ in $\mathsf{E}$.
Furthermore, in both cases the restricted cotorsion pair $(\mathsf{E}\cap\mathsf{F},\,\mathsf{E}\cap\mathsf{C})$ in $\mathsf{E}$ is hereditary whenever the cotorsion pair $(\mathsf{F},\mathsf{C})$ in $\mathsf{K}$ is hereditary.
Proof.
This is [33, Lemmas 1.5 and 1.6]. ∎
An exact category $\mathsf{E}$ is said to have exact directed colimits if all directed colimits exist in $\mathsf{E}$ and preserve (admissible) short exact sequences. The following important proposition goes back to [45, proofs of Lemma 5.2 and Proposition 5.3], [18, Propositions 3.1 and 3.2], and [2, proofs of Lemma 4.6 and Theorem 4.7].
Proposition 7.4.
Let $\mathsf{E}$ be an exact category with exact directed colimits and $\mathsf{A}\subset\mathsf{E}$ be a full subcategory closed under extensions and cokernels of admissible monomorphisms in $\mathsf{E}$. Then $\mathsf{A}$ is closed under transfinitely iterated extensions if and only if $\mathsf{A}$ is closed under directed colimits in $\mathsf{E}$.
Proof.
This is [42, Proposition 8.1]. ∎
In the following lemma, the class of objects $\mathsf{F}={}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$ is viewed as a full subcategory of $\mathsf{K}$ endowed with the exact structure inherited from $\mathsf{K}$.
Lemma 7.5.
Let $\mathsf{K}$ be an exact category and $\mathsf{C}\subset\mathsf{K}$ be a full subcategory. Put $\mathsf{F}={}^{\perp_{1}}\mathsf{C}\subset\mathsf{K}$. Let $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be a complex in $\mathsf{C}$. In this context:
(a) For any (termwise admissible) short exact sequence $0\longrightarrow F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow G^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow H^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ of complexes in $\mathsf{F}$, the short sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(H^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(G^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(F^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$ is exact.
(b) For any absolutely acyclic complex $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ in $\mathsf{F}$, any morphism of complexes $A^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
This is a generalization of the dual version of Lemma 2.5(a). Part (a) is obvious; and to prove part (b), one can observe that acyclicity of the complex of abelian groups $\operatorname{Hom}_{\mathsf{K}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(A^{\text{\smaller\smaller$\scriptstyle\bullet$}},C^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ follows from part (a). ∎
In the next lemma, both the classes $\mathsf{F}$ and $\mathsf{C}\subset\mathsf{K}$ are viewed as full subcategories endowed with the exact structures inherited from $\mathsf{K}$.
Lemma 7.6.
Let $\mathsf{K}$ be an exact category and $(\mathsf{F},\mathsf{C})$ be a cotorsion pair in $\mathsf{K}$. Let $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be an acyclic complex in the exact category $\mathsf{F}$ and $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be an acyclic complex in the exact category $\mathsf{C}$. Then, viewing both $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and $C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ as complexes in $\mathsf{K}$, any morphism of complexes $F^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow C^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero.
Proof.
This is [15, Lemma 3.9] or [29, Lemma B.1.8]. ∎
8. Cotorsion Pair with Complexes of Cotorsion Contramodules
The aim of this section is to prove a contramodule version of the theorem of Bazzoni, Cortés-Izurdiaga, and Estrada [2, Theorem 5.3]. The argument in [2] is based on cotorsion periodicity together with [15, Corollary 4.10]. The proof of the latter uses considerations of deconstructibility in the category of complexes.
Such deconstructibility results have been established in full generality for Grothendieck abelian categories [44, Section 4], which does not cover our context. (Indeed, the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ are not exact, while $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ is not an abelian category.) Our proof in this section does not use that much knowledge about deconstructibility, but is based on accessibility considerations instead.
We start with the following proposition going back to [7, Lemma 1 and Proposition 2]. The definition of a deconstructible class was given in Section 2.
Proposition 8.1.